Jason and Marie both travel 900 miles. Jason drove at an average speed of 10 miles per hour faster than Marie and made the trip in one hour less time. Find the average speeds of both cars.
Jason's speed = ?mph
Marie's speed = ?mph

Answer by stanbon(75887)

Jason and Marie both travel 900 miles. Jason drove at an average speed of 10 miles per hour faster than Marie and made the trip in one hour less time. Find the average speeds of both cars.
----
Jason DATA:
distance = 900 miles ; rate = x+10 mph ; time = 900/(x+10) hrs
-----
Marie DATA:
distance = 900 miles ; rate = x mph ; time = 900/x hrs
------
Equation:
time - time = 1 hr
900/x - 900/(x+10) = 1 hr
900x+9000 - 900x = x(x+10)
x^2+10x-9000 = 0
(x+100)(x-90) = 0
Positive solution:
x = 90 mph (Marie's speed)
x+10 = 100 mph (Jason's speed)
===================================
Cheers,
Stan H.
============



Jason's speed = ?mph
Marie's speed = ?mph